Multi-wavelength observations of compact solar flares G. Del Zanna H.E. Mason DAMTP, University of Cambridge, UK A. Berlicki, B. Schmieder (Observatoire de Paris,Meudon)

Contents

1 Introduction 1 1.1 paper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Overview 1

3 The M1 flare 1 3.1 Zoom in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

4 Post-flare CDS raster 15:43–17:07 UT 1

5 Conclusions 1

1. Introduction

– The SOHO/Coronal Diagnostic Spectrometer provides essential spectroscopic information to understand the characteristics and dynamics of solar coronal features.

– We give a description of the main observational features and what types of measurements (densities, temperatures, flows) are obtained.

– These provide important observational constraints for flare modeling.

– CDS observations of a compact M1 flare that occurred on October 22 2002, also observed from the ground in chromospheric lines and by RHESSI. The flare occurred in a region of strong and mixed magnetic polarity, was compact and lasted only a short time.

1.1. paper

2. Overview

THEMIS observed continuously between 15:41–17:26 UT.

TRACE and VTT/MSDP observations were performed during the whole flare. VTT/MSDP observations in Hα 9 scans between 15:30 and 16:17 UT TRACE high cadence (40 s) in the 195 Å band between 15:29 and 15:41 UT

2 Fig. 1. THEMIS/MSDP in Na I D_{1} (15:41–16:29 UT) and line-of-sight magnetic field deduced from I and V Stokes overlaid with Hα contours observed by VTT/MSDP at 15:30:01 UT.

3. The M1 flare 3.1. Zoom in

4. Post-flare CDS raster 15:43–17:07 UT

5. Conclusions

– Strong blue-shifts in lines emitted in the 1-8 MK range are observed at the bases of hot Fe XIX structures.

Fig. 2. RHESSI counts rates (the 12–25 keV band are multiplied by a factor of 10).

– The hot structures are also clearly visible in the TRACE 195 Å band

– The hot structures match the thermal emission as seen by RHESSI

Acknowledgements Support from PPARC is acknowledged.

Fig. 3. Top: TRACE 195 Å + MDI contours Bottom: CDS FOV - TRACE 195 Å + contours of Fe XIX intensity 7 8 9